OpenStudy (anonymous):

At a music store, 8 CDs cost $32. How much will 15 CDs cost at this store?

OpenStudy (anonymous):

15*4=answer

OpenStudy (anonymous):

32/8=4 per CD 4*15 = answer

OpenStudy (fozia):

60 dollers

OpenStudy (anonymous):

32/8=4 15*4=60

OpenStudy (anonymous):

Thank you everyone.